# Demystifying Deep Search: A Holistic Evaluation with Hint-free Multi-Hop Questions and Factorised Metrics

### v2

@@ 4.2.1 Understanding Model Failure Modes Through Metric Patterns
Table 2: Emergent model profiles from metric interplay analysis.
@@ 4.2.2 Knowledge Degradation in Synthesis
From table 1, models with lower Forget and Lead-astray generally exhibit higher Knowledge Utilization, which in turn coincides with higher Generation Score and Pass@1. For instance, Grok-4 (Forget 23.65%, Lead-astray 27.70%) attains the highest Knowledge Utilization F1 (56.19%), the highest Generation Score (34.71%), and 50.5% Pass@1. Similarly, o3-Pro (Forget 21.83%, Lead-astray 25.35%) reaches 49.40% Knowledge Utilization, 20.86% Generation Score, and the best Pass@1 (56.0%). GPT-5 shows a comparable pattern with very low Forget (17.72%) and strong Knowledge Utilization (49.58%), alongside 23.21% Generation Score and 50.5% Pass@1. In contrast, when Forget is high, Knowledge Utilization collapses and downstream metrics follow suit: GPT-OSS-120B records the lowest Knowledge Utilization (10.73%) with Forget at 100.00% (Lead-astray 0.00%), yielding only 2.75% Generation Score and 24.0% Pass@1; Qwen3-30B-Thinking has Forget 79.65% (Lead-astray 16.81%), with 13.16% Knowledge Utilization, 7.25% Generation Score, and 12.5% Pass@1. Similar degradations appear for Gemini-2.5-Flash-Think (Forget 57.63%, Knowledge Utilization 16.35%) and Tongyi-DeepResearch-30B (Forget 43.93%, Knowledge Utilization 15.69%).
@@ 4.2.3 Robustness to Test-Time Scaling
Figure 2: Scaling under test-time scaling (TTS).
@@
In fig. 2, we observe two main trends. For Claude-Opus-4.1, enlarging the context window from 8K to 32K tokens brings negligible gains: Generation Score plateaus at about 34%, Pass@1 at about 50%, and Search Score increases by less than 1%. For EvidenceLoop, expanding the controller from breadth=1, iteration=2 to breadth=3, iteration=2 raises Search Score slightly (45% $\rightarrow$ 46%, +1%), leaves Generation Score unchanged at 21%, and improves Pass@1 from 49% to 56% (+7%). These results indicate that our benchmark is robust to naïve test-time scaling. Neither larger context budgets nor shallow exploration suffice to “hack” WebDetective; achieving further gains requires genuine advances in model reasoning and knowledge utilisation.
@@ 5 Conclusion
We introduce WebDetective, a benchmark for evaluating web agents on hint-free multi-hop deep search within a controlled Wikipedia sandbox. Unlike prior datasets that embed reasoning paths (PH) or entity fingerprints (SH), our design enforces autonomous discovery of reasoning chains while enabling fine-grained attribution of failure modes. Evaluation of 25 state-of-the-art models reveals consistent weaknesses: systems often retrieve sufficient evidence but fail to utilise it effectively, and appropriate refusals remain nearly absent. Our proposed agentic workflow EvidenceLoop demonstrates that explicit verification and systematic evidence tracking can partially close this gap, underscoring that performance cannot be trivially improved by test-time scaling alone.
@@ Appendix A Formal Metrics Definition
$\text{SearchScore}=\text{KnowledgeScore}+\frac{|\mathcal{C}|}{N}$ (2)
where $\mathcal{C}=\{d\in\mathcal{D}:\text{correct}(d)\land\text{hops}(d)\leq\text{hops}_{\text{GT}}(d)\land\text{KS}(d)=0\}$ represents instances where the model:
  • Uses no more search hops than the ground truth path length
This metric recognizes that effective multi-hop reasoning isn’t solely about following predetermined paths, but about efficiently discovering and leveraging available information—whether through complete evidence chains or intelligent combination of partial search with existing knowledge. The hop constraint ensures models aren’t simply performing exhaustive search, but are discovering meaningful connections that enable efficient reasoning.
@@ Appendix C Dataset Statistics
Appendix D The EvidenceLoop framework
@@
D.1 Core Architecture: Iterative Refinement with Fallback
@@
D.2 Evidence Memory System
@@
The EID system serves multiple critical functions in our framework. First, during extraction and aggregation between iterations, the extraction and aggregation agent produces summaries that preserve EID references alongside extracted facts—for example, “Kane has brother Chad [EID-042], Chad’s stepmother is Nicole [EID-089]”. This allows subsequent solver agents to receive concise, actionable summaries while retaining the ability to retrieve full evidence on demand through the retrieve action as an external tool, which takes an EID and returns the complete original content from memory. Second, these EIDs enable systematic verification (detailed in Section D.3), where verification agents can trace claims back to their original sources and validate reasoning chains against the actual evidence.
@@
D.3 Verification: Ensuring Evidence-Grounded Reasoning
@@
Appendix E Failure Case Studies
@@
Appendix F Related Work
F.1 Multi-Hop Question Answering Benchmarks
@@
Multi-hop QA benchmarks evaluate models’ ability to compose information across multiple reasoning steps. Early datasets like HotpotQA Yang et al. (2018b) and WikiHop Welbl et al. (2018) established foundational evaluation frameworks but suffer from systematic biases. Recent benchmarks have expanded coverage: FanOutQA Zhu et al. (2024) addresses multi-document reasoning, MINTQA He et al. (2024) targets long-tail knowledge with 28K+ questions, and MEQA Anonymous (2024a) focuses on event-centric reasoning chains. However, these benchmarks embed hints that fundamentally alter the reasoning task.
@@
We identify two categories of hints prevalent in existing benchmarks. Path-hinting occurs when questions linguistically encode reasoning chains (e.g., “What dance academy did the starring actress from The Glory of Tang Dynasty graduate from?”), reducing the task to executing pre-specified steps. Specification-hinting provides excessive constraints that make answers discoverable through constraint satisfaction rather than reasoning (e.g., combining “East German team,” “founded 1966,” “player born in 90s”). Unlike MuSiQue Trivedi et al. (2022) or 2WikiMultiHopQA Ho et al. (2020), which contain implicit structural hints, WebDetective introduces genuinely hint-free questions requiring autonomous reasoning path discovery.
F.2 Retrieval-Augmented Generation and Agents
@@
The evolution from static RAG pipelines to agentic architectures represents a fundamental shift in how LLMs interact with external knowledge Singh et al. (2025); Ehtesham et al. (2025). While traditional RAG systems like TRACE Fang et al. (2024) achieve improvements through knowledge-grounded reasoning chains, they operate within predetermined patterns. Agentic RAG systems employ adaptive strategies: Adaptive-RAG Jeong et al. (2024) adjusts retrieval depth based on question complexity, while graph-based approaches like GNN-Ret Li et al. (2024b) and HopRAG Liu et al. (2025) leverage graph neural networks for multi-hop reasoning, achieving  10% accuracy improvements on benchmarks like 2WikiMQA.
@@
Recent advances in 2025 emphasize diverse reasoning paths. DP-CoT Li et al. (2024a) addresses single-path limitations through passage-level and sentence-level evidence generation. However, our evaluation reveals these advances fail to overcome hint-free challenges: median Generation Scores of 20% across tested models indicate current architectures cannot effectively discover reasoning chains without linguistic scaffolding.
F.3 Evaluation Frameworks
@@
Traditional metrics like exact match and F1 scores collapse diverse failure modes into single values, obscuring why models fail Kwiatkowski et al. (2019); Petroni et al. (2021). Recent frameworks attempt more nuanced evaluation: RAGAS Shahul et al. (2024) provides reference-free RAG metrics, while RAGTruth Niu et al. (2024) enables hallucination analysis. For agents, AgentBench Liu et al. (2023) evaluates across eight environments, tau-bench AI (2024) addresses multi-turn interactions, and TheAgentCompany Anonymous (2024b) introduces workplace tasks with simulated colleagues.
@@
Web-based benchmarks have evolved significantly. WebArena Zhou et al. (2023) provides realistic web environments requiring long-horizon planning but lacks controlled evaluation for precise failure attribution. SWE-bench Jimenez et al. (2024) evaluates code generation on GitHub issues, with SWE-bench Verified OpenAI (2024) addressing underspecified problems. While these benchmarks test complex capabilities, they don’t address the specific challenge of verifying multi-hop reasoning paths.

### v3

@@ 4.2.1 Understanding Model Failure Modes Through Metric Patterns
Table 3: Emergent model profiles from metric interplay analysis. Models cluster into distinct behavioural profiles based on their knowledge sufficiency, refusal behaviour, and information synthesis, revealing characteristic strengths and failure modes across the spectrum of frontier systems.
@@ 4.2.2 Knowledge Degradation in Synthesis
From Table 2, models with lower Forget and Lead-astray rates consistently achieve stronger Knowledge Utilization, which in turn drives higher Generation Score and Pass@1. High performers such as Grok-4 (Forget 23.65%, Lead-astray 27.70%) and o3-Pro (Forget 21.83%, Lead-astray 25.35%) maintain the best Knowledge Utilization (56.19% and 49.40%), translating into the top Generation Scores (34.71% and 20.86%) and competitive Pass@1 (50.5% and 56.0%). GPT-5 shows a similar profile with very low Forget (17.72%) and strong Knowledge Utilization (49.58%). In contrast, models with severe knowledge degradation see downstream metrics collapse. GPT-OSS-120B, with Forget at 100%, attains only 10.73% Knowledge Utilization and 2.75% Generation Score; Qwen3-30B-Thinking (Forget 79.65%) similarly drops to 13.16% Knowledge Utilization and 7.25% Generation. Mid-tier models such as Gemini-2.5-Flash-Think and Tongyi-DeepResearch-30B follow the same pattern. Overall, the results show that stable evidence retention—not raw retrieval—is the key determinant of deep-search reasoning performance.
@@ 4.2.3 Robustness to Test-Time Scaling
Figure 2: Test-time scaling (TTS) on WebDetective. Left: Increasing Claude-Opus-4.1’s context length boosts Search Score and Pass@1 at shorter ranges, but both plateau beyond 32k tokens, while Generation Score shows only modest gains. Right: EvidenceLoop remains stable across breadth–iteration settings, with Pass@1 improving under moderate configurations (e.g., 1–2, 3–2). Generation Score changes little, highlighting synthesis—not search—as the main bottleneck under TTS.
@@
In Figure 2, we observe two main trends. For Claude-Opus-4.1, enlarging the context window from 8K to 32K tokens brings negligible gains: Generation Score plateaus at about 34%, Pass@1 at about 50%, and Search Score increases by less than 1%. For EvidenceLoop, expanding the controller from breadth=1, iteration=2 to breadth=3, iteration=2 raises Search Score slightly (45% $\rightarrow$ 46%, +1%), leaves Generation Score unchanged at 21%, and improves Pass@1 from 49% to 56% (+7%). These results indicate that our benchmark is robust to naïve test-time scaling. Neither larger context budgets nor shallow exploration suffice to “hack” WebDetective; achieving further gains requires genuine advances in model reasoning and knowledge utilisation.
@@ 5 Conclusion
We introduce WebDetective, a benchmark for evaluating web agents on hint-free multi-hop deep search within a controlled Wikipedia sandbox. Unlike prior datasets that embed reasoning paths (PH) or entity fingerprints (SH), our design enforces autonomous discovery of reasoning chains while enabling fine-grained attribution of failure modes. Crucially, while the sandbox eliminates corpus-based shortcuts, it accommodates multiple valid reasoning strategies—agents may combine partial search with parametric knowledge, and our metrics credit any path achieving knowledge sufficiency. Evaluation of 25 state-of-the-art models reveals consistent weaknesses: systems often retrieve sufficient evidence but fail to utilise it effectively, and appropriate refusals remain nearly absent. Our proposed agentic workflow EvidenceLoop, designed to target the specific failure modes our diagnostics reveal, demonstrates that explicit verification and systematic evidence tracking can partially close this gap—improving Knowledge Utilization by 53% over its base model—underscoring that performance cannot be trivially improved by test-time scaling alone. The co-design methodology generalises beyond Wikipedia to any domain with factual content, entity graphs, and masking capability, establishing WebDetective as a framework for developing genuinely autonomous reasoning systems.
@@ Appendix C Dataset Statistics
Domain Coverage. The dataset spans diverse topical domains to ensure broad evaluation coverage, as shown in Table 4. This distribution ensures WebDetective is not skewed toward any single relation type (such as familial ties) but instead covers reasoning patterns essential for comprehensive evaluation.
Table 4: Domain distribution in the WebDetective benchmark.
Domain Percentage
Film, Television & Theatre 20.0%
History, Politics & Military 18.0%
Geography, Locations & Landmarks 16.0%
Music & Audio Production 13.5%
Sports & Competitions 11.5%
Biography, Genealogy & Education 8.5%
Literature, Comics & Pop Culture 6.5%
Science, Technology & Transport 6.0%
Appendix D Statistical Considerations
Quality-Focused Construction Pipeline. The relatively modest dataset size (200 questions) is a direct consequence of our strict construction and filtering pipeline. For each candidate question-answer pair $(q,a^{*})$ with evidence chain $\mathcal{E}=\{e_{1},...,e_{n}\}$, we enforce three conditions: (1) Parametric Inaccessibility: the verifier LM cannot produce $a^{*}$ from $q$ alone; (2) Evidence Sufficiency: $\text{LM}(q,\mathcal{E})=a^{*}$; and (3) Evidence Necessity: for every $i\in\{1,...,n\}$, we probe with the incomplete set $(q,\mathcal{E}\setminus\{e_{i}\})$ and verify failure. The majority of automatically generated candidates are discarded during this rigorous process, followed by manual verification.
Confidence Intervals. Our binomial confidence analysis shows that with $n=200$, we achieve a 95% confidence interval of $\pm 6.9$% (assuming 50% pass rate). While this is sufficient for identifying the large-scale failure patterns our benchmark targets—the gap between Knowledge Score (79%) and Generation Score (23%) for top models, the near-universal failure of refusal capabilities ($30\%$ Good Refusal F1), and the synthesis bottleneck (Knowledge Utilization peaks at 56%)—these are substantial effects that do not require larger samples to detect reliably.
Diagnostic Purpose. Our primary aim with WebDetective is to provide a diagnostic benchmark that exposes fundamental failure modes rather than ranking systems with marginal differences. The patterns we identify are consistent across all 25 evaluated systems, demonstrating that even with our 2-3 hop concentration (86% of questions), the benchmark remains far from saturated: the highest Pass@1 is only 56% (o3-Pro), and state-of-the-art agents frequently fail to synthesize 2-3 pieces of evidence without hints.
Appendix E The EvidenceLoop framework
@@
E.1 Core Architecture: Iterative Refinement with Fallback
@@
E.2 Evidence Memory System
@@
The EID system serves multiple critical functions in our framework. First, during extraction and aggregation between iterations, the extraction and aggregation agent produces summaries that preserve EID references alongside extracted facts—for example, “Kane has brother Chad [EID-042], Chad’s stepmother is Nicole [EID-089]”. This allows subsequent solver agents to receive concise, actionable summaries while retaining the ability to retrieve full evidence on demand through the retrieve action as an external tool, which takes an EID and returns the complete original content from memory. Second, these EIDs enable systematic verification (detailed in Section E.3), where verification agents can trace claims back to their original sources and validate reasoning chains against the actual evidence.
@@
E.3 Verification: Ensuring Evidence-Grounded Reasoning
@@
E.4 Ablations on Architecture Design
Table 5: Ablation study on EvidenceLoop. Search Score and Knowledge Utilization F1 are reported, with emphasis on the memory and verification modules, identified earlier as key failure points.
Model Variant Search Score (%) Knowledge Util. F1 (%)
EvidenceLoop (DeepSeek-R1-0528) 71.50 25.00
    - w/o Memory 63.00 24.38
    - w/o Verification 68.60 23.46
To quantify the contributions of the memory and verification components, we conduct ablation studies using DeepSeek-R1-0528 as the base model. As shown in Table 5, memory-driven working-context structuring is the main force behind alleviating knowledge-related deficiencies, while verification plays a complementary role by strengthening precision and ensuring robust utilization of retrieved knowledge.
Impact of Memory.
Removing the structured evidence buffer (w/o Memory) leads to the most substantial performance drop, with the Search Score decreasing by 8.5% (71.5% $\to$ 63.0%). This underscores that memory-driven working-context structuring is the primary contributor to mitigating knowledge-related shortages and enabling effective knowledge discovery.
Impact of Verification.
Eliminating the explicit verification loop (w/o Verification) results in a moderate but consistent reduction in both Search Score (-2.9%) and Knowledge Utilization F1 (-1.54%). This indicates that verification provides valuable additional support, improving retrieval quality and enhancing the stability of information synthesis.
E.5 Generalization Across Base Models
To assess whether the benefits of EvidenceLoop rely on specific model characteristics or generalize across different backbones, we extended our evaluation to two distinct architectures: DeepSeek-R1-0528 and GLM-4.5-Air. This analysis specifically tests the framework’s robustness against context degradation and limited search coverage inherent in standard ReAct-style baselines.
As detailed in Table 6, EvidenceLoop delivers consistent improvements in knowledge discovery and synthesis across both architectures. Specifically, we observe:
  • Search Coverage: Search Scores improve by +3.0+3.0 to +6.0+6.0 points, indicating a more effective exploration of the information space.
  • Knowledge Utilization: The structured evidence buffer and verification loop yield substantial gains in utilization, with relative improvements in F1 Score ranging from $10\%$ to $53\%$.
  • Task Performance: These gains translate directly to downstream accuracy, boosting Pass@1 by +4.5+4.5 to +8.0+8.0 points absolute ($24\%$ to $39\%$ relative).
Most notably, the DeepSeek-R1-0528 backbone achieves a $53\%$ relative improvement in Knowledge Utilization F1 (16.36% $\to$ 25.0%). This result confirms that EvidenceLoop effectively mitigates the synthesis bottleneck identified by WebDetective, transforming retrieved raw context into actionable evidence through its iterative explore-verify mechanism.
Table 6: Performance comparison across different base models. EvidenceLoop consistently outperforms the standard baselines on DeepSeek-R1-0528 and GLM-4.5-Air, particularly in Knowledge Utilization and overall Pass@1 accuracy. Relative improvements are indicated in parentheses.
Model & Method Knol. Suff Search Sc. Knol. Util F1 Pass@1
Backbone: DeepSeek-R1-0528
  Baseline 61.0 65.5 16.36 20.5
  EvidenceLoop 68.0 71.5 25.00 28.5
$\Delta$ +7.0 +6.0 +53% (rel.) +39% (rel.)
Backbone: GLM-4.5-Air
  Baseline 55.5 60.5 17.97 19.0
  EvidenceLoop 62.0 63.5 19.70 23.5
$\Delta$ +6.5 +3.0 +10% (rel.) +24% (rel.)
Appendix F Failure Case Studies
@@
Appendix G Related Work
G.1 Multi-Hop Question Answering Benchmarks
@@
Multi-hop QA benchmarks evaluate models’ ability to compose information across multiple reasoning steps. Early datasets like HotpotQA yang2018hotpotqa and WikiHop welbl2018constructing established foundational evaluation frameworks but suffer from systematic biases. Recent benchmarks have expanded coverage: FanOutQA zhu2024fanoutqa addresses multi-document reasoning, MINTQA he2024mintqa targets long-tail knowledge with 28K+ questions, and MEQA neurips2024meqa focuses on event-centric reasoning chains. However, these benchmarks embed hints that fundamentally alter the reasoning task.
@@
We identify two categories of hints prevalent in existing benchmarks. Path-hinting occurs when questions linguistically encode reasoning chains (e.g., “What dance academy did the starring actress from The Glory of Tang Dynasty graduate from?”), reducing the task to executing pre-specified steps. Specification-hinting provides excessive constraints that make answers discoverable through constraint satisfaction rather than reasoning (e.g., combining “East German team,” “founded 1966,” “player born in 90s”). Unlike MuSiQue trivedi2022musique or 2WikiMultiHopQA ho20202wikimultihopqa, which contain implicit structural hints, WebDetective introduces genuinely hint-free questions requiring autonomous reasoning path discovery.
G.2 Retrieval-Augmented Generation and Agents
@@
The evolution from static RAG pipelines to agentic architectures represents a fundamental shift in how LLMs interact with external knowledge singh2025agentic; ehtesham2025systematic. While traditional RAG systems like TRACE fang2024trace achieve improvements through knowledge-grounded reasoning chains, they operate within predetermined patterns. Agentic RAG systems employ adaptive strategies: Adaptive-RAG jeong2024adaptive adjusts retrieval depth based on question complexity, while graph-based approaches like GNN-Ret li2024gnnret and HopRAG liu2025hoprag leverage graph neural networks for multi-hop reasoning, achieving  10% accuracy improvements on benchmarks like 2WikiMQA.
@@
Recent advances in 2025 emphasize diverse reasoning paths. DP-CoT li2024dpCoT addresses single-path limitations through passage-level and sentence-level evidence generation. However, our evaluation reveals these advances fail to overcome hint-free challenges: median Generation Scores of 20% across tested models indicate current architectures cannot effectively discover reasoning chains without linguistic scaffolding.
G.3 Evaluation Frameworks
@@
Traditional metrics like exact match and F1 scores collapse diverse failure modes into single values, obscuring why models fail kwiatkowski2019natural; petroni2021kilt. Recent frameworks attempt more nuanced evaluation: RAGAS shahul2024ragas provides reference-free RAG metrics, while RAGTruth niu2024ragtruth enables hallucination analysis. For agents, AgentBench liu2023agentbench evaluates across eight environments, tau-bench sierra2024taubench addresses multi-turn interactions, and TheAgentCompany theagentcompany2024 introduces workplace tasks with simulated colleagues.
@@
Web-based benchmarks have evolved significantly. WebArena zhou2023webarena provides realistic web environments requiring long-horizon planning but lacks controlled evaluation for precise failure attribution. SWE-bench jimenez2024swebench evaluates code generation on GitHub issues, with SWE-bench Verified openai2024swebenchverified addressing underspecified problems. While these benchmarks test complex capabilities, they don’t address the specific challenge of verifying multi-hop reasoning paths.
@@
Appendix H LLMs Usage
LLMs were used to polish the writing.